# BayesFLo: Bayesian fault localization of complex software systems

### v2

@@ Abstract
Software testing is essential for the reliable development of complex software systems. A key step in software testing is fault localization, which uses test data to pinpoint failure-inducing combinations for further diagnosis. Existing fault localization methods have two key limitations: they (i) do not incorporate domain and/or structural knowledge from test engineers, and (ii) do not provide a probabilistic assessment of risk for potential root causes. Such methods can thus fail to confidently whittle down the combinatorial number of potential root causes in complex systems, resulting in prohibitively high testing costs. To address this, we propose a novel Bayesian fault localization framework called BayesFLo, which leverages a flexible Bayesian model for identifying potential root causes with probabilistic uncertainty. Using a carefully-specified prior on root cause probabilities, BayesFLo permits the integration of domain and structural knowledge via the principles of combination hierarchy and heredity, which capture the expected structure of failure-inducing combinations. We then develop new algorithms for efficient computation of posterior root cause probabilities, leveraging recent tools from integer programming and graph representations. Finally, we demonstrate the effectiveness of BayesFLo over existing methods in two fault localization case studies on the Traffic Alert and Collision Avoidance System and the JMP Easy DOE platform.
@@ 1 Introduction
Software testing – the process of executing a program with the intent of finding errors (Myers et al.,, 2004) – is an essential step in the development of robust software applications. Such testing aims to reveal and subsequently fix as many bugs (i.e., faults) as possible prior to the deployment of a software application, thus greatly reducing the likelihood of encountering failures for the end-user. This is crucial in an era where nearly all facets of daily life involve human interaction with software applications. There are, however, two fundamental challenges. First, each software test can be time-consuming to perform. This involves not only running the software system itself, which can be intensive for modern scientific applications (Ji et al., 2024a, ; Ji et al., 2024b, ), but also determining whether the software deviates from its expected behavior. The latter, known as the “oracle problem” (Barr et al.,, 2014), typically requires an independent assessment of numerical accuracy of software outputs (Lekivetz and Morgan,, 2021), which can be very costly. Second, the number of test cases required for thorough software examination can easily be overwhelming. As “bugs [tend to] lurk in corners and congregate at boundaries” (Beizer,, 2003), software testing typically focuses on boundary values and the combinations of inputs, which can grow rapidly. For practical applications, it is thus wholly infeasible to exhaustively test all input combinations (Kumar,, 2019), which can easily require billions of test cases.
@@
These two fundamental challenges open up a world of exciting new statistical directions for this application of rising importance. Such directions can roughly be categorized into two topics. The first is a careful design of test cases to perform with the joint goals of identifying failure settings and diagnosing its underlying root causes. Statistically, this can be viewed as an experimental design problem for software fault diagnosis. There has been notable work on this front. Early work on one-factor-at-a-time designs (Frey et al.,, 2003) explored the problem of unit testing (Runeson,, 2006; Khorikov,, 2020), which focuses on testing small functional code units, e.g., a single function or chunk of code. However, such testing cannot identify bugs that congregate at the interaction of such units, which often arise in large software applications. Another approach is pairwise testing (Bach and Schroeder,, 2004), which examines all pairwise combinations of inputs; test case generation for this setting has been explored in Tai and Lei, (2002). A more general approach is combinatorial testing (Nie and Leung, 2011b, ), which investigates combinations involving more than two inputs. For combinatorial testing, the design of choice is a covering array (CA; Colbourn,, 2004; Dalal and Mallows,, 1998); such designs aim to represent (or “cover”) each combination of inputs (up to a specified order) at least once. CAs are thus ideal for detecting failures from limited test runs; we will discuss CAs in greater detail in Section 2.
@@
With the initial test cases conducted and failures detected, the second direction is fault localization (Wong et al.,, 2023): the use of this test data to pinpoint root causes. This is a highly challenging problem due to the overwhelming number of scenarios to consider for potential root causes. To see why, consider a software application with $I=10$ input factors each with two levels. For this system, there are a total of $\sum_{i=1}^{10}\binom{10}{i}2^{i}=59,048$ input combinations that might be potential root causes; these consist of all single factors, two-factor combinations, and so on. Since each combination is either a root cause or not, this results in a whopping $2^{59048}$ different scenarios to consider for potential root causes! Fault localization requires gauging which of these many scenarios is likely given test set outcomes, which is clearly a computationally intensive task (Wong et al.,, 2023), even for systems with a moderate number of inputs $I$ and few failed test cases. Furthermore, the debugging cost for investigating each such scenario can be considerably high (Boehm,, 1984; Humphrey,, 1995). For example, for critical flight software systems at the Jet Propulsion Laboratory (Bush,, 1990), each debugging activity can cost over thousands of dollars to identify and fix faults.
@@
Due to this combinatorial explosion of potential root causes and high debugging costs, there is thus an urgent need for fault localization methods that, based on initial test case data, can identify a small number of suspicious input combinations for further investigation. On this front, Nie and Leung, 2011a proposed a minimal failure-causing schema and used it to narrow down the search range for potential root causes and to guide subsequent test case generation. Niu et al., (2013) proposed a notion of tuple relationship tree for visualizing the relationships among all input combinations. Such a tree is utilized to eliminate “healthy” combinations and to propose subsequent test cases for further system examination. More recently, Ghandehari, (2016) and Ghandehari et al., (2018) introduced a two-phase approach for finding faulty statements in a software system. Such approaches have been integrated, either in full or in part, within the Covering Array Analysis module in the statistical software package JMP (henceforce called JMP; Jones and Sall,, 2011).
The above body of work, however, has two key limitations that hamper its effectiveness in practical applications. First, they are not probabilistic in nature, and thus offer little insight on the probability of a combination being a root cause given test outcomes. Such probabilities are needed for confident fault localization; they provide a principled measure of risk on the “suspiciousness” of a root cause. A probabilistic approach can further accelerate fault localization by disentangling the many tied combinations given limited test runs; more on this later. Second, existing methods do not permit the integration of domain and/or structural knowledge from test engineers within a probabilistic framework. Here, domain knowledge refers to prior information that may raise or reduce the suspiciousness of certain input factors; this may come, e.g., from historical data, expert judgment or specific characteristics of the system. Structural knowledge, in contrast, captures expected fault behavior from an understanding of the software’s architecture and how components interact (Lekivetz and Morgan,, 2018, 2021). Existing methods that do not integrate such knowledge may fail to whittle down the combinatorial number of potential root causes, which leads to costly and/or delayed fault diagnosis, as we shall see in our applications later.
@@
To address these challenges, we propose a new Bayesian Fault LOcalization (BayesFLo) framework. The main workhorse of BayesFLo is a new Bayesian model on root cause indicators over all possible input combinations, which facilitates easy integration of domain knowledge from test engineers. This model further embeds the desirable principles of combination hierarchy and heredity (Lekivetz and Morgan,, 2021), which capture structural knowledge on software root causes. We show that the integration of such principles, derived from the well-known principles of effect hierarchy and heredity (Wu and Hamada,, 2009) for analyzing experimental data, can accelerate the identification of root causes from limited test cases. A critical challenge for Bayesian computation is the sheer number of considered combinations, which renders the computation of posterior root cause probabilities to be infeasible without careful manipulation. We thus develop a new algorithm for efficient computation of such posterior probabilities, leveraging recent tools from integer programming and graph representations. Finally, we demonstrate the practical advantages of BayesFLo over existing methods in two fault localization case studies, the first for the Traffic alert and Collision Avoidance System (TCAS; Bradley,, 1992) and the other for the Easy DOE platform in JMP (Jones and Sall,, 2011).
@@
The paper is organized as follows. Section 2 outlines our motivating application on fault localization of TCAS, as well as limitations of the state-of-the-art. Section 3 presents the BayesFLo model, and describes the combination hierarchy and heredity principles embedded in its prior specification. Section 4 proposes a novel algorithm for efficient computation of posterior root cause probabilities for potential root cause combinations. Section 5 explores the application of BayesFLo in two practical case studies on the fault localization of TCAS and the JMP Easy DOE platform. Section 6 concludes the paper.
@@ 2.1 Background & Challenges
Our motivating application is in the fault localization of the Traffic alert and Collision Avoidance System (TCAS; Harman,, 1989; Williamson and Spencer,, 1989), which is jointly developed by the Federal Aviation Administration (FAA) and the aviation industry. TCAS is an aircraft collision avoidance system, designed to prevent potential mid-air collisions between aircraft (FAA,, 2011). Figure 1 visualizes its operation. TCAS monitors the airspace around an aircraft, and determines a protection volume around the aircraft based on its speed and trajectory. When two aircraft come too close either horizontally or vertically, TCAS alerts the pilots by issuing a resolution advisory. TCAS is currently mandated for U.S. commercial aircraft, and is also installed on many helicopters and turbine-powered general aviation aircraft (FAA,, 2011). Given the catastrophic implications of aircraft collisions, e.g., from recent events (Davies and Yousif,, 2025), it is paramount to mitigate such risks by ensuring the reliability of TCAS systems via comprehensive testing under different scenarios.
Figure 1: Visualizing the protection volume module in the TCAS software system (FAA,, 2011).
@@ 12 Climb_Inhibit 0, 1
Testing the TCAS software system has attracted much attention within the software engineering community, and there has been fruitful work on designing test suites for detecting defects in the system (Griesmayer et al.,, 2007; Ghandehari et al.,, 2013). In what follows, we consider a specific TCAS program from the TCAS software system – the Siemens test suite in Griesmayer et al., (2007) – for fault localization. This program takes in $I=12$ input factors (summarized in Table 1 with its corresponding levels), with each factor parameterizing certain characteristics of the aircraft system. Of these $I=12$ factors, nine pertain to the primary aircraft, and the remaining three (with prefix Other in Table 1) are for an incoming aircraft that has entered its protection volume (Xie et al.,, 2013). The latter consists of inputs 6, 10 and 11, which reflect the altitude, resolution advisory and TCAS capability of the incoming aircraft. The former consists of remaining inputs that reflect the operating conditions of the primary aircraft, e.g., its current altitude, altitude rate-of-change, and separation altitudes with the incoming aircraft.
@@
A key challenge here is the expensive testing cost for each run of TCAS. Recall that a “failure” in a software program is when its actual behavior deviate from its expected behavior. Here, the assessment of the latter requires the construction of eleven data collection sites over the U.S. (Bradley,, 1992) and the collection of ground radar data (spanning hundreds of hours) for generating encounter databases. Logic test simulations are performed on thousands of such encounters (Bradley,, 1992), and physical experiments are then conducted to validate its real-world performance. Coupled with the many potential root cause combinations from the $I=12$ input factors, such testing costs make the exhaustive testing of TCAS over all combinations wholly impractical.
@@
An effective solution is to first construct a small designed test set of input settings. An appealing design employed in the literature is a covering array (CA; Colbourn,, 2004; Lekivetz and Morgan,, 2021), which covers each input combination (up to a certain order) at least once. Formally, take an array with dimensions $M\times I$, and suppose its $i$-th column takes $J_{i}$ distinct levels for some integer $J_{i}\geq 2$. Then this array is a CA of strength $s$, if for any choice of $s$ columns, each level combination involving such columns occurs at least once. For software testing, this CA can be used to design a test set with $M$ runs and $I$ inputs, where the $i$-th input has $J_{i}$ levels; the levels in its $m$-th row specify the input settings for the $m$-th test run. Table 2 shows a strength-2 CA for our TCAS case study. Such a design achieves the desired coverage condition with $M=17$ runs, which greatly reduces the number of validation runs on the expensive software system. With this CA, all single inputs and two-factor input combinations are investigated in at least one test run; if there is a bug within such combinations, we would observe a corresponding failure in a test run.
@@
The next step is to use this test run data to identify potential root causes, i.e., the problem of fault localization. The last column in Table 2 shows the outcomes (0 for a passed case, 1 for a failed case) for each test run. A key challenge here is that there are more than $\sum_{i=1}^{12}\binom{12}{i}2^{i}>500,000$ input combinations (e.g., the two-factor combination Cur_Vertical_Sep = 299 and High_Confidence = 0) that may be failure-inducing root causes. Since each combination is either a root cause or not, there are thus a whopping $2^{500,000}$ different root cause scenarios to consider for fault localization! The goal is to parse through these many scenarios to find which are likely given the observed test results. With the costly nature of debugging, this set of likely scenarios should further consist of only a few combinations, to facilitate cost-efficient downstream diagnosis. Another challenge is the need for assessing confidence that an identified suspicious combination is indeed a root cause. This provides test engineers a principled way to decide which combinations are likely root causes that need to be further investigated, and which are likely not root causes and can be safely ignored. Such uncertainty quantification is critical for trustworthy fault diagnosis (Zhou et al.,, 2023). We inspect next existing methods through the lens of these two challenges.
@@
Table 2: The $M=17$-run CA design and test outcomes for our motivating TCAS case study. Here, an outcome of 0 indicates a passed test case, with 1 indicating a failed test case.

### v3

@@ Abstract
Software testing is essential for the reliable development of complex software systems. A key step in software testing is fault localization, which uses test data to pinpoint failure-inducing combinations for further diagnosis. Existing fault localization methods have two key limitations: they (i) largely do not incorporate domain and/or structural knowledge from test engineers, and (ii) do not provide a probabilistic assessment of risk for potential root causes. Such methods can thus fail to confidently whittle down the combinatorial number of potential root causes in complex systems, resulting in prohibitively high debugging costs. To address this, we propose a novel Bayesian fault localization framework called BayesFLo, which leverages a flexible Bayesian model for identifying potential root causes with probabilistic uncertainties. Using a carefully-specified prior on root cause probabilities, BayesFLo permits the integration of domain and structural knowledge via the principles of combination hierarchy and heredity, which capture the expected structure of failure-inducing combinations. We then develop new algorithms for efficient computation of posterior root cause probabilities, leveraging recent tools from integer programming and graph representations. Finally, we demonstrate the effectiveness of BayesFLo over existing methods in two fault localization case studies, the first on the Traffic Alert and Collision Avoidance System for aircraft collision avoidance, and the second on the Vulnerable Road User protection tests for safe autonomous driving.
@@ 1 Introduction
Software testing – the process of executing a program with the intent of finding errors (Myers et al., 2004) – is an essential step in the development of robust software applications. Such testing aims to reveal and subsequently fix as many bugs (i.e., faults) as possible prior to the deployment of a software application, thus greatly reducing the likelihood of encountering failures for the end-user. This is crucial in an era where nearly all facets of daily life involve human interaction with software applications. There are, however, two fundamental challenges. First, each software test can be time-consuming to perform. This involves not only running the software system itself, which can be intensive for modern scientific applications (Ji et al., 2024a, b), but also determining whether the software deviates from its expected behavior. The latter, known as the “oracle problem” (Barr et al., 2014), typically requires an independent assessment of numerical accuracy of software outputs (Lekivetz and Morgan, 2021), which can be very costly. Second, the number of test cases required for thorough software examination can easily be overwhelming. As “bugs [tend to] lurk in corners and congregate at boundaries” (Beizer, 2003), software testing typically focuses on boundary values and the combinations of inputs, which can grow rapidly. For practical applications, it is thus wholly infeasible to exhaustively test all input combinations (Kumar, 2019), which can easily require billions of test cases.
@@
These two fundamental challenges open up exciting new statistical directions for this application of rising importance. Such directions can roughly be categorized into two topics. The first is a careful design of test cases with the joint goals of identifying failure settings and diagnosing their root causes. Statistically, this can be viewed as an experimental design problem for software fault diagnosis. There has been notable work on this front. Early work on one-factor-at-a-time designs (Frey et al., 2003) explored the problem of unit testing (Runeson, 2006; Khorikov, 2020), which focuses on testing small functional code units, e.g., a single function or chunk of code. However, such testing cannot identify bugs that congregate at the interaction of such units, which often arise in large software applications. Another approach is pairwise testing (Bach and Schroeder, 2004), which examines all pairwise combinations of inputs; test case generation for this setting has been explored in Tai and Lei (2002). A more general approach is combinatorial testing (Nie and Leung, 2011a), which investigates combinations involving more than two inputs. For combinatorial testing, the design of choice is a covering array (CA; Colbourn, 2004), which aims to represent (or “cover”) each combination of inputs (up to a specified order) at least once. CAs enable the detection of failures from limited test runs; see Section 2 for further discussion.
@@
With the initial test cases conducted and failures detected, the second direction is fault localization (Wong et al., 2023): the use of this test data to pinpoint root causes. This is a highly challenging problem due to the overwhelming number of scenarios to consider for potential root causes. To see why, consider a software application with $I=10$ input factors each with two levels. For this system, there are a total of $\sum_{i=1}^{10}\binom{10}{i}2^{i}=59,048$ input combinations that might be potential root causes; these consist of all single factors, 2-factor combinations, and so on. Since each combination is either a root cause or not, this results in $2^{59048}$ different scenarios to consider for potential root causes! Fault localization requires gauging which of these many scenarios are likely given test outcomes, which is a computationally intensive task (Wong et al., 2023), even for systems with a moderate number of inputs and few failed test cases. Furthermore, the debugging cost for investigating each such scenario can be considerably high (Boehm, 1984; Humphrey, 1995). For example, for critical flight software systems at the Jet Propulsion Laboratory (Bush, 1990), each debugging activity can cost thousands of dollars to identify and fix faults.
@@
Due to this combinatorial explosion of potential root causes and high debugging costs, there is an urgent need for fault localization methods that, based on initial test case data, can identify a small number of suspicious input combinations (i.e., potential root causes) for further investigation. On this front, Nie and Leung (2011b) proposed a minimal failure-causing schema, which narrows down the search range for potential root causes and guides subsequent test case generation. Niu et al. (2013) proposed a notion of tuple relationship tree for visualizing the relationships among all input combinations. Such a tree is utilized to eliminate “healthy” combinations and to propose subsequent test cases for further system examination. Ghandehari (2016) and Ghandehari et al. (2020) introduced a two-phase approach for finding faulty statements in a software system. Lekivetz and Morgan (2018) proposed a deterministic ranking procedure that incorporates a domain-knowledge-guided weighting scheme.
Despite this literature, there remains a need for a principled statistical framework for fault localization. Such a framework, implemented in a Bayesian fashion, provides three practical advantages for efficient and confident fault localization. First, it can accelerate the identification of a few key suspicious combinations from the combinatorially many potential root causes using test data, enabling efficient fault diagnosis with a small number of test runs. In particular, it permits the principled comparison of potential root causes with differing numbers of inputs, which addresses a key limitation of existing methods (see Section 2). Second, a Bayesian framework facilitates the integration of domain and/or structural knowledge from test engineers via its prior specification. Here, domain knowledge refers to prior information from historical data, expert judgment or system characteristics that raise or reduce the suspiciousness of certain input factors. Structural knowledge captures fault behavior from an understanding of the software’s architecture and how components interact (Lekivetz and Morgan, 2018, 2021). A careful integration of such knowledge can accelerate the identification of root causes when faced with an overwhelming list of potential causes. Finally, a Bayesian framework enables principled risk quantification, by providing for each suspicious combination the posterior probability that it may be a root cause given test outcomes. Such probabilities support confident fault localization by guiding the required debugging effort. If only several combinations have high posterior probabilities, then debugging costs are expected to be low; if many combinations have non-negligible probabilities, then a massive debugging effort may be needed to fully investigate such combinations. Recent work (Landsberg et al., 2016; Landsberg and Barr, 2018) points to an increasing need for probabilistic metrics for risk-aware fault localization to gain deeper insights on software failure behavior.
@@
We propose a new Bayesian Fault LOcalization (BayesFLo) framework that effectively achieves these three advantages. The main workhorse of BayesFLo is a new Bayesian model on root cause indicators over all possible input combinations, which facilitates easy integration of domain knowledge from test engineers. This model further embeds the desirable principles of combination hierarchy and heredity (Lekivetz and Morgan, 2021), which capture structural knowledge on software root causes. We show that the integration of such principles, derived from the well-known principles of effect hierarchy and heredity (Wu and Hamada, 2009) for analyzing factorial experiments, can accelerate the identification of root causes from limited test runs. A key challenge for Bayesian computation is the sheer number of considered combinations, which renders the computation of posterior root cause probabilities to be infeasible without careful manipulation. We develop a new algorithm for efficient computation of such posterior probabilities, leveraging recent tools from integer programming and graph representations. Finally, we demonstrate the practical advantages of BayesFLo over existing methods in two fault localization case studies, on the Traffic Alert and Collision Avoidance System (TCAS; Bradley, 1992) for aircraft collision avoidance, and the Vulnerable Road User (VRU; Gladisch et al., 2020) protection tests for safe autonomous driving.
@@
The paper is organized as follows. Section 2 outlines our motivating application on fault localization of TCAS, as well as limitations of the state-of-the-art. Section 3 presents the BayesFLo model, and describes the combination hierarchy and heredity principles embedded in its prior specification. Section 4 proposes a novel algorithm for efficient computation of posterior root cause probabilities for potential root cause combinations. Section 5 explores the application of BayesFLo in two practical case studies on the fault localization of the TCAS system and the VRU protection tests. Section 6 concludes the paper.
@@ 2.1 Background & Challenges
Our motivating application is in the fault localization of the Traffic Alert and Collision Avoidance System (TCAS; Harman, 1989; Williamson and Spencer, 1989), which is jointly developed by the Federal Aviation Administration (FAA) and the aviation industry. TCAS is designed to prevent potential mid-air collisions between aircraft (FAA, 2011). Figure 1 visualizes its operation. TCAS monitors the airspace around an aircraft, and determines a protection volume around the aircraft based on its speed and trajectory. When two aircraft come too close, TCAS alerts the pilots by issuing a resolution advisory. TCAS is currently mandated for U.S. commercial aircraft, and is also installed on many helicopters and turbine-powered general aviation aircraft (FAA, 2011). Given the catastrophic implications of aircraft collisions, e.g., from recent events (Davies and Yousif, 2025), it is important to mitigate such risks by ensuring the reliability of TCAS systems via comprehensive testing under different scenarios.
Figure 1: Visualizing the protection volume module in the TCAS software system (FAA, 2011).
@@ 12 Climb_Inhibit 0, 1
Testing the TCAS software system has attracted much attention within the software engineering community, and there has been fruitful work on designing test suites for detecting defects in the system (Griesmayer et al., 2007; Ghandehari et al., 2013). In what follows, we consider a specific TCAS program from the TCAS software system – the Siemens test suite in Griesmayer et al. (2007) – for fault localization. This program takes in $I=12$ input factors (summarized in Table 1), with each factor parameterizing certain characteristics of the aircraft system. Of these $I=12$ factors, nine pertain to the primary aircraft, and the remaining three (with prefix Other in Table 1) are for an incoming aircraft that has entered its protection volume (Xie et al., 2013). The latter consists of inputs 6, 10 and 11, which reflect the altitude, resolution advisory and TCAS capability of the incoming aircraft. The former consists of remaining inputs that reflect the operating conditions of the primary aircraft, e.g., its current altitude, altitude rate-of-change, and separation altitudes with the incoming aircraft.
@@
A fundamental challenge here is the expensive testing cost for each run of TCAS. Recall that a “failure” in a software program is when its actual behavior deviates from its expected behavior. Here, the assessment of the latter requires the construction of eleven data collection sites over the U.S. (Bradley, 1992) and the collection of ground radar data (spanning hundreds of hours) for generating encounter databases. Logic test simulations are performed on thousands of such encounters (Bradley, 1992), and physical experiments are then conducted to validate its real-world performance. Such testing costs make the exhaustive testing of TCAS over all combinations of the $I=12$ input factors wholly impractical.
@@
An effective solution is to first construct a small designed set of inputs for testing. A reasonable design is a covering array (CA; Colbourn, 2004), which covers each input combination (up to a certain order) at least once. Formally, take an array with dimensions $M\times I$, and suppose its $i$-th column takes $J_{i}$ distinct levels for some integer $J_{i}\geq 2$. Then this array is a CA of strength $s$, if for any choice of $s$ columns, each level combination involving such columns occurs at least once. For software testing, this CA can be used to design a test set with $M$ runs and $I$ inputs, where the $i$-th input has $J_{i}$ levels; the levels in its $m$-th row specify the input settings for the $m$-th test run. Table 2 shows a strength-2 CA for our TCAS case study. This design achieves the desired coverage with $M=19$ runs, which reduces the number of runs on the expensive software system. Using this CA, all single inputs and 2-factor input combinations are investigated in at least one test run; if there is a bug within such combinations, we would observe a corresponding failure in a test run.
@@
The next step is to use these test outcomes to identify and rank potential root causes, i.e., fault localization. The last column in Table 2 shows the outcomes (0 for a passed case, 1 for a failed case) for each test run. Here, a key challenge is that there are more than $\sum_{i=1}^{12}\binom{12}{i}2^{i}>500,000$ input combinations (e.g., the 2-factor combination Cur_Vertical_Sep = 299 and High_Confidence = 0) that may be failure-inducing root causes. Since each combination is either a root cause or not, there are thus $2^{500,000}$ different root cause scenarios to consider for fault localization! The goal is to parse through these many scenarios to find which are likely given the observed test results. With the costly nature of debugging, this set of likely scenarios should further consist of only a few combinations, to facilitate cost-efficient downstream diagnosis. Another challenge is the need to assess, for a suspicious combination, the probability this is indeed a root cause given test outcomes. This provides test engineers a principled way to gauge which combinations are likely root causes that need to be further investigated, and which can be confidently ignored. This probabilistic assessment of risk is desirable for trustworthy fault diagnosis (Landsberg and Barr, 2018; Zhou et al., 2023). We inspect next existing methods through the lens of these two challenges.
@@
18 299 0 1 1 600 1 1 300 840 0 1 1 0
19 601 1 0 1 600 1 3 0 640 2 1 0 0
Table 2: The $M=19$-run CA design and test outcomes for our motivating TCAS case study. Here, an outcome of 0 indicates a passed test case, with 1 indicating a failed test case.